# Robust estimation of boundary using doubly stochastic scaling of Gaussian kernel

### v2

@@ 6. Proofs
Proof of Lemma 6.1. Let $(\partial_{i})_{1}^{d}$ be an orthonormal basis of $T_{\mathbf{x}}\mathcal{M}$. Then there exist $a^{i}\in\mathbb{R}$ and $b^{i}:\mathcal{M}\mapsto\mathbb{R}$ for all $i\in[1,d]$ such that $\mathbf{v}_{\mathbf{x}}=a^{i}\partial_{i}$ (using Einstein notation) and $\mathcal{P}_{\mathbf{x}}(\mathbf{y})=b^{i}(\mathbf{y})\partial_{i}$. Moreover, $b^{i}(\mathbf{x})=0$ for all $i\in[1,d]$,
@@
Proof of Theorem 2.1. For brevity, we use the following shorthand notation for the proof: $\rho\equiv\rho_{h}(\mathbf{x})$, $q\equiv q(\mathbf{x})$, $H\equiv H(\mathbf{x})$, $\omega\equiv\omega(\mathbf{x})$, $A\equiv A_{h}(\mathbf{x})$, $B\equiv B_{h}(\mathbf{x})$, $m_{i}\equiv m_{h}^{(i)}(\mathbf{x})$ for $i=1,2$ and $3$.
@@
[…]
Using $\partial_{\mathbf{v}_{\mathbf{x}}}k_{h}(\mathbf{x},\mathbf{y})=2k_{h}(\mathbf{x},\mathbf{y})\mathbf{v}_{\mathbf{x}}^{T}\mathcal{P}_{\mathbf{x}}(\mathbf{y})/h^{2}$ and Eq. 8 followed by Lemmas 2.1 and 6.1,
(50) $\pi^{d/2}\frac{\partial_{\mathbf{v}_{\mathbf{x}}}\rho}{\rho}+\frac{2m_{1}\rho^{2}q\ \mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}}}{h}+2\widetilde{m}_{2}\rho\left(q\partial_{\mathbf{v}_{\mathbf{x}}}\rho+\rho\partial_{\mathbf{v}_{\mathbf{x}}}q+\left(\frac{m_{2}}{\widetilde{m}_{2}}-1\right)\partial_{\bm{\eta}_{\mathbf{x}}}(\rho q)\mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}}\right)+\mathcal{O}(h)=0.$
By substituting $\mathbf{v}_{\mathbf{x}}=\bm{\eta}_{\mathbf{x}}$ in Eq. 50 and simplifying, we obtain
(51) $(\pi^{d/2}+2m_{2}\rho^{2}q)h\partial_{\bm{\eta}_{\mathbf{x}}}\rho+2\rho^{3}(m_{1}q+hm_{2}\partial_{\bm{\eta}_{\mathbf{x}}}q)+\mathcal{O}(h^{2})=0.$
Substituting $\mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}}=0$ in Eq. 50,
@@
Simplifying the above and substituting the value of $\widetilde{m}_{2}=\pi^{d/2}/2$ as defined in Eq. 18, we obtain
@@
(56) $=\rho q\left\{(m_{0}+hm_{1}\omega)\rho+hm_{1}\partial_{\bm{\eta}_{\mathbf{x}}}\rho\right\}+\mathcal{O}(h^{2})$
@@
(57) $hm_{1}\rho q\partial_{\bm{\eta}_{\mathbf{x}}}\rho=\pi^{d/2}-(m_{0}+hm_{1}\omega)\rho^{2}q+\mathcal{O}(h^{2}).$
Multiplying Eq. 51 by $m_{1}\rho q$ and substituting Eq. 57 results in
@@
where, using the definition of $\omega$ in the statement of the theorem,
(59) $A$ $=2\left(m_{1}^{2}+hm_{1}m_{2}\frac{\partial_{\bm{\eta}_{\mathbf{x}}}q}{q}-m_{2}(m_{0}+hm_{1}\omega)\right)$
(60) $=2\left(m_{1}^{2}-m_{2}\left(m_{0}+\frac{h(d-1)m_{1}H}{2}\right)\right)$
@@
(61) $B=\pi^{d/2}(m_{0}+hm_{1}\omega-2m_{2})=\pi^{d/2}m_{1}\left(h\omega-\frac{2b_{\mathbf{x}}}{h}\right).$
Proof of Corollary 2.1. The proof follows from Eq. 57 by using the fact $b_{\mathbf{x}}=\infty$ for every point on a closed manifold, and as a result of substituting it in Eq. 15 and 16, we obtain $m^{(0)}_{h}(\mathbf{x})=\pi^{d/2}$ and $m^{(1)}_{h}(\mathbf{x})=0$.
Proof of Corollary 2.2. We adopt the same notation as in the proof of Theorem 2.1. Due to the assumption $H=0$, we have
@@
therefore $\rho=C_{-}+\mathcal{O}(h^{2})$. Since $\mathcal{M}$ is compact therefore $C_{-}$ is bounded and thus $\rho=C_{-}(1+\mathcal{O}(h^{2}))$. Finally, if $\partial_{\bm{\eta}_{\mathbf{x}}}q$ is also zero then $\omega=0$ by definition and the result follows.
Proof of Corollary 2.3. Substituting $b_{\mathbf{x}}=0$, we obtain $m^{(0)}_{h}(\mathbf{x})=\frac{\pi^{d/2}}{2}$, $m^{(1)}_{h}(\mathbf{x})=-\frac{\pi^{(d-1)/2}}{2}$ and $m^{(2)}_{h}(\mathbf{x})=\frac{\pi^{d/2}}{4}$.
Proof of Corollary 2.4. This follows trivially from Theorem 2.1 by taking away the terms involving $hH(\mathbf{x})$ and $h\omega(\mathbf{x})$.
Proof of Theorem 3.1. Using assumption (A6) followed by (A7), we have
@@
[…]
@@
From here on, for brevity, we use the shorthand notation as in the proof of Theorem 2.1. Due to Lemma 2.1 and the product rule for second order derivatives,
[…]
Denote by $\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}=\nabla(\rho q)-\partial_{\bm{\eta}_{\mathbf{x}}}(\rho q)\bm{\eta}_{\mathbf{x}}$. Using Theorem 2.1, the r.h.s reduces to,
(68) $hm_{1}\rho q\bm{\eta}_{\mathbf{x}}+h^{2}(\widetilde{m}_{2}\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}+m_{2}\partial_{\bm{\eta}_{\mathbf{x}}}(\rho q)\bm{\eta}_{\mathbf{x}})+\mathcal{O}\left(h^{3}\right)$
(69) $=hm_{1}\rho q\bm{\eta}_{\mathbf{x}}+h^{2}(\widetilde{m}_{2}\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}+m_{2}(\rho\partial_{\bm{\eta}_{\mathbf{x}}}q+q\partial_{\bm{\eta}_{\mathbf{x}}}\rho)\bm{\eta}_{\mathbf{x}})+\mathcal{O}\left(h^{3}\right)$
@@
(70) $=h\left(m_{1}\rho q-\frac{2m_{1}m_{2}\rho^{3}q^{2}}{\pi^{d/2}+2m_{2}\rho^{2}q}\right)\bm{\eta}_{\mathbf{x}}+h^{2}\left(\widetilde{m}_{2}\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}+m_{2}\left(\rho\partial_{\bm{\eta}_{\mathbf{x}}}q-\frac{2m_{2}\rho^{3}q\partial_{\bm{\eta}_{\mathbf{x}}}q}{\pi^{d/2}+2m_{2}\rho^{2}q}\right)\bm{\eta}_{\mathbf{x}}\right)+\mathcal{O}\left(h^{3}\right)$
@@
(71) $=h\left(\frac{\pi^{d/2}m_{1}\rho q}{\pi^{d/2}+2m_{2}\rho^{2}q}\right)\bm{\eta}_{\mathbf{x}}+h^{2}\left(\widetilde{m}_{2}\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}+\frac{\pi^{d/2}m_{2}\rho\partial_{\bm{\eta}_{\mathbf{x}}}q}{\pi^{d/2}+2m_{2}\rho^{2}q}\bm{\eta}_{\mathbf{x}}\right)+\mathcal{O}\left(h^{3}\right)$
From Theorem 2.1, if $\mathbf{v}_{\mathbf{x}}\perp\bm{\eta}_{\mathbf{x}}$ then
(73) $\partial_{\mathbf{v}_{\mathbf{x}}}(\rho q)=(\rho\partial_{\mathbf{v}_{\mathbf{x}}}q+q\partial_{\mathbf{v}_{\mathbf{x}}}\rho)=\left(\rho\partial_{\mathbf{v}_{\mathbf{x}}}q-\frac{\rho^{3}q\partial_{\mathbf{v}_{\mathbf{x}}}q}{1+\rho^{2}q}\right)+\mathcal{O}(h)=\frac{\rho\partial_{\mathbf{v}_{\mathbf{x}}}q}{1+\rho^{2}q}+\mathcal{O}(h).$
Using the above equation and Corollary 2.4 (which also shows $\rho^{2}q=\mathcal{O}(1)$), the r.h.s. reduces to,
@@
(74) $h\left(\frac{\pi^{d/2}m_{1}\rho q}{\pi^{d/2}+2m_{2}\rho^{2}q}\right)\bm{\eta}_{\mathbf{x}}+\frac{h^{2}\pi^{d/2}\rho}{2}\left(\frac{\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}q}{1+\rho^{2}q}\bm{\eta}_{\mathbf{x}_{\perp}}+\frac{2m_{2}\partial_{\bm{\eta}_{\mathbf{x}}}q}{\pi^{d/2}+2m_{2}\rho^{2}q}\bm{\eta}_{\mathbf{x}}\right)+\mathcal{O}\left(h^{3}\right)$
(75) $=\frac{h}{\rho}\left(\frac{\pi^{d/2}m_{1}\zeta}{\pi^{d/2}+2m_{2}\zeta}\right)\bm{\eta}_{\mathbf{x}}+\mathcal{O}\left(h^{2}\right).$

### v3

@@ Appendix A Proofs
Proof.
Let $(\partial_{i})_{1}^{d}$ be an orthonormal basis of $T_{\mathbf{x}}\mathcal{M}$. Then there exist $a^{i}\in\mathbb{R}$ and $b^{i}:\mathcal{M}\mapsto\mathbb{R}$ for all $i\in[1,d]$ such that $\mathbf{v}_{\mathbf{x}}=a^{i}\partial_{i}$ (using Einstein notation) and $\mathcal{P}_{\mathbf{x}}(\mathbf{y})=b^{i}(\mathbf{y})\partial_{i}$. Moreover, $b^{i}(\mathbf{x})=0$ for all $i\in[1,d]$,
@@
Proof of Theorem 3.1.
For brevity, we use the following shorthand notation for the proof: $\rho\equiv\rho_{h}(\mathbf{x})$, $q\equiv q(\mathbf{x})$, $H\equiv H(\mathbf{x})$, $\kappa_{\pm}\equiv\kappa_{\pm}(\mathbf{x})$, $A\equiv A_{h}(\mathbf{x})$, $B\equiv B_{h}(\mathbf{x})$, $m_{i}\equiv m_{h}^{(i)}(\mathbf{x})$ for $i=1,2$ and $3$, and $\widetilde{m}_{2}=\frac{\pi^{d/2}}{2}$.
@@
(53) $\frac{1}{h^{d}}\int_{\mathcal{M}}(k_{h}(\mathbf{x},\mathbf{y})\partial_{\mathbf{v}_{\mathbf{x}}}\rho_{h}(\mathbf{x})+\rho_{h}(\mathbf{x})\ \partial_{\mathbf{v}_{\mathbf{x}}}k_{h}(\mathbf{x},\mathbf{y}))\rho_{h}(\mathbf{y})q(\mathbf{y})dV(\mathbf{y})=0.$
For the first term, using Lemma 3.1 and Eq. 9, we obtain:
(54) $\frac{1}{h^{d}}\int_{\mathcal{M}}k_{h}(\mathbf{x},\mathbf{y})\partial_{\mathbf{v}_{\mathbf{x}}}\rho_{h}(\mathbf{x})\rho_{h}(\mathbf{y})q(\mathbf{y})dV(\mathbf{y})$
(55) $\hskip 28.45274pt=\pi^{d/2}\frac{\partial_{\mathbf{v}_{\mathbf{x}}}\rho_{h}(\mathbf{x})}{\rho_{h}(\mathbf{x})}\frac{1}{\pi^{d/2}h^{d}}\int_{\mathcal{M}}k_{h}(\mathbf{x},\mathbf{y})\rho_{h}(\mathbf{x})\rho_{h}(\mathbf{y})q(\mathbf{y})dV(\mathbf{y})$
(56) $\hskip 28.45274pt=\pi^{d/2}\frac{\partial_{\mathbf{v}_{\mathbf{x}}}\rho_{h}(\mathbf{x})}{\rho_{h}(\mathbf{x})}.$
Then, for the second term, using $\partial_{\mathbf{v}_{\mathbf{x}}}k_{h}(\mathbf{x},\mathbf{y})=2k_{h}(\mathbf{x},\mathbf{y})\mathbf{v}_{\mathbf{x}}^{T}\mathcal{P}_{\mathbf{x}}(\mathbf{y})/h^{2}$ followed by Lemmas 3.1 and A.1 and the following facts
(57) $\Delta(f\mathbf{v}_{\mathbf{x}}^{T}\mathcal{P}_{\mathbf{x}})(\mathbf{x})$ $=\mathbf{v}_{\mathbf{x}}^{T}\mathcal{P}_{\mathbf{x}}(\mathbf{x})\Delta f(\mathbf{x})+f(\mathbf{x})\Delta(\mathbf{v}_{\mathbf{x}}^{T}\mathcal{P}_{\mathbf{x}})(\mathbf{x})+2\partial_{\mathbf{v}_{\mathbf{x}}}f(\mathbf{x})$
(58) $=0+0+2\partial_{\mathbf{v}_{\mathbf{x}}}f(\mathbf{x})$
(59) $\partial_{\bm{\eta}_{\mathbf{x}}}^{2}(f\mathbf{v}_{\mathbf{x}}^{T}\mathcal{P}_{\mathbf{x}})(\mathbf{x})$ $=0+0+2\partial_{\bm{\eta}_{\mathbf{x}}}f(\mathbf{x})\mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}}$
(60) $\partial_{\bm{\eta}_{\mathbf{x}}}(f\mathbf{v}_{\mathbf{x}}^{T}\mathcal{P}_{\mathbf{x}})(\mathbf{x})$ $=f(\mathbf{x})\mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}},$
we obtain the following equations:
(61) $\frac{2\rho_{h}(\mathbf{x})}{h^{d+2}}\int_{\mathcal{M}}k_{h}(\mathbf{x},\mathbf{y})\mathbf{v}_{\mathbf{x}}^{T}\mathcal{P}_{\mathbf{x}}(\mathbf{y})\rho_{h}(\mathbf{y})q(\mathbf{y})dV(\mathbf{y})$
(62) $\hskip 14.22636pt=\frac{2m_{h}^{(1)}(\mathbf{x})\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})\mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}}}{h}+\rho_{h}(\mathbf{x})\left\{2\widetilde{m}_{2}\partial_{\mathbf{v}_{\mathbf{x}}}(\rho_{h}q)(\mathbf{x})+\right$.
(63) $\hskip 35.56593pt\left.2(m_{h}^{(2)}(\mathbf{x})-\widetilde{m}_{2})\partial_{\bm{\eta}_{\mathbf{x}}}(\rho_{h}q)(\mathbf{x})\mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}}-(d-1)m_{h}^{(2)}(\mathbf{x})H(\mathbf{x})\rho_{h}(\mathbf{x})q(\mathbf{x})\mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}}\right\}+\mathcal{O}(h)$
By substituting $\mathbf{v}_{\mathbf{x}}=\bm{\eta}_{\mathbf{x}}$ in Eq. 56 and Eq. 63, and equating the sum to zero, we obtain
(64) $\pi^{d/2}\frac{\partial_{\bm{\eta}_{\mathbf{x}}}\rho}{\rho}+\frac{2m_{1}\rho^{2}q}{h}+\rho\left\{2m_{2}\partial_{\bm{\eta}_{\mathbf{x}}}(\rho q)-(d-1)m_{2}H\rho q\right\}+\mathcal{O}(h)$ $=0$
(65) $(\pi^{d/2}+2m_{2}\rho^{2}q)h\partial_{\bm{\eta}_{\mathbf{x}}}\rho+2\rho^{3}(m_{1}q+hm_{2}\partial_{\bm{\eta}_{\mathbf{x}}}q)-(d-1)hm_{2}H\rho^{3}q+\mathcal{O}(h^{2})$ $=0$
(66) $(\pi^{d/2}+2m_{2}\rho^{2}q)h\partial_{\bm{\eta}_{\mathbf{x}}}\rho+2\rho^{3}q(m_{1}+hm_{2}\kappa_{-})+\mathcal{O}(h^{2})$ $=0$
Similarly, substituting $\mathbf{v}_{\mathbf{x}}^{T}\bm{\eta}_{\mathbf{x}}=0$ in Eq. 56 and Eq. 63, and equating the sum to zero, we obtain
@@
Simplifying the above and substituting the value of $\widetilde{m}_{2}=\pi^{d/2}/2$, we obtain
@@
(71) $=\rho q\left\{(m_{0}+hm_{1}\kappa_{+})\rho+hm_{1}\partial_{\bm{\eta}_{\mathbf{x}}}\rho\right\}+\mathcal{O}(h^{2})$
@@
(72) $hm_{1}\rho q\partial_{\bm{\eta}_{\mathbf{x}}}\rho=\pi^{d/2}-(m_{0}+hm_{1}\kappa_{+})\rho^{2}q+\mathcal{O}(h^{2}).$
Multiplying Eq. 66 by $m_{1}\rho q$,
(73) $(\pi^{d/2}+2m_{2}\rho^{2}q)hm_{1}\rho q\partial_{\bm{\eta}_{\mathbf{x}}}\rho+2m_{1}\rho^{4}q^{2}(m_{1}+hm_{2}\kappa_{-})+\mathcal{O}(h^{2})=0$
and then substituting Eq. 72 results in
@@
where, using the definition of $\kappa_{+}$ in the statement of the theorem,
(75) $A$ $=2\left(m_{1}^{2}+hm_{1}m_{2}\kappa_{-}-m_{2}(m_{0}+hm_{1}\kappa_{+})\right)$
(76) $=2\left(m_{1}^{2}-m_{2}\left(m_{0}+h(d-1)m_{1}H\right)\right)$
@@
(77) $B=\pi^{d/2}(m_{0}+hm_{1}\kappa_{+}-2m_{2})=\pi^{d/2}m_{1}\left(h\kappa_{+}-\frac{2b_{\mathbf{x}}}{h}\right).$
Proof of Corollary 3.1.
The proof follows from Eq. 72 by using the fact $b_{\mathbf{x}}=\infty$ for every point on a closed manifold, and as a result of substituting it in Eq. 19 and 20, we obtain $m^{(0)}_{h}(\mathbf{x})=\pi^{d/2}$ and $m^{(1)}_{h}(\mathbf{x})=0$. ∎
Proof of Corollary 3.2.
We adopt the same notation as in the proof of Theorem 3.1. Due to the assumption $H=0$, we have
@@
therefore $\rho^{2}q=C_{-}+\mathcal{O}(h^{2})$. Since $\mathcal{M}$ is compact therefore $C_{-}$ is bounded and thus $\rho^{2}q=C_{-}(1+\mathcal{O}(h^{2}))$. Finally, if $\partial_{\bm{\eta}_{\mathbf{x}}}q$ is also zero then $\kappa_{+}=0$ by definition and the result follows. ∎
Proof of Corollary 3.3.
Setting $b_{\mathbf{x}}=0$, we obtain $m^{(0)}_{h}(\mathbf{x})=\frac{\pi^{d/2}}{2}$, $m^{(1)}_{h}(\mathbf{x})=-\frac{\pi^{(d-1)/2}}{2}$ and $m^{(2)}_{h}(\mathbf{x})=\frac{\pi^{d/2}}{4}$. ∎
Proof of Corollary 3.4.
This follows trivially from Theorem 3.1 by taking away the terms involving $hH(\mathbf{x})$ and $h\kappa_{+}(\mathbf{x})$ and solving the quadratic in $\rho_{h}^{2}q$. ∎
Proof of Theorem 4.1.
Using assumption (A6) followed by (A7), we have
@@
(82) $\bm{\nu}_{i}=\frac{\rho_{h}(\mathbf{x}_{i})}{\pi^{d/2}h^{d}}\ \bm{\mathcal{U}}_{i}^{T}\left(\int k_{h}(\mathbf{x}_{i},\mathbf{y})\rho_{h}(\mathbf{y})(\mathbf{y}-\mathbf{x}_{i})q(\mathbf{y})dV(\mathbf{y})\right)+\mathcal{O}^{(h)}_{m,n}(g(m,n)).$
@@
From here on, for brevity, we use the shorthand notation as in the proof of Theorem 3.1 and use $\mathbf{x}$ in place of $\mathbf{x}_{i}$. Due to Lemma 3.1 and the product rule for second order derivatives,
@@
(83) $\frac{1}{h^{d}}\int k_{h}(\mathbf{x},\mathbf{y})\rho_{h}(\mathbf{y})(\mathbf{y}-\mathbf{x})q(\mathbf{y})dV(\mathbf{y})=hm_{1}\rho q\bm{\eta}_{\mathbf{x}}+h^{2}\left(\widetilde{m}_{2}\nabla(\rho q)+\left(m_{2}-\widetilde{m}_{2}\right)\partial_{\bm{\eta}_{\mathbf{x}}}(\rho q)\bm{\eta}_{\mathbf{x}}\right)+\mathcal{O}\left(h^{2}\right)\bm{\eta}_{\mathbf{x}}$
and from Theorem 3.1 we have,
(84) $h\partial_{\bm{\eta}_{\mathbf{x}}}\rho=-\frac{2m_{1}\rho^{3}q}{\pi^{d/2}+2m_{2}\rho^{2}q}+\mathcal{O}(h)$
and, if $\mathbf{v}_{\mathbf{x}}\perp\bm{\eta}_{\mathbf{x}}$ then
(85) $\partial_{\mathbf{v}_{\mathbf{x}}}(\rho q)=(\rho\partial_{\mathbf{v}_{\mathbf{x}}}q+q\partial_{\mathbf{v}_{\mathbf{x}}}\rho)=\left(\rho\partial_{\mathbf{v}_{\mathbf{x}}}q-\frac{\rho^{3}q\partial_{\mathbf{v}_{\mathbf{x}}}q}{1+\rho^{2}q}\right)+\mathcal{O}(h)=\frac{\rho\partial_{\mathbf{v}_{\mathbf{x}}}q}{1+\rho^{2}q}+\mathcal{O}(h),$
where the constants in $\mathcal{O}\left(h^{2}\right)$ and $\mathcal{O}\left(h\right)$, respectively, can be bounded uniformly over $\mathbf{x}\in\mathcal{M}$ (this holds throughout the proof). Let $\bm{\eta}_{\mathbf{x}_{\perp}}$ represent the subspace orthogonal to $\bm{\eta}_{\mathbf{x}}$, then $\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}=\nabla(\rho q)-\partial_{\bm{\eta}_{\mathbf{x}}}(\rho q)\bm{\eta}_{\mathbf{x}}$. Using the above equations, we obtain
(86) $hm_{1}\rho q\bm{\eta}_{\mathbf{x}}+h^{2}(\widetilde{m}_{2}\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}+m_{2}\partial_{\bm{\eta}_{\mathbf{x}}}(\rho q)\bm{\eta}_{\mathbf{x}})+\mathcal{O}\left(h^{2}\right)\bm{\eta}_{\mathbf{x}}$
(87) $=hm_{1}\rho q\bm{\eta}_{\mathbf{x}}+h^{2}(\widetilde{m}_{2}\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}+m_{2}(\rho\partial_{\bm{\eta}_{\mathbf{x}}}q+q\partial_{\bm{\eta}_{\mathbf{x}}}\rho)\bm{\eta}_{\mathbf{x}})+\mathcal{O}\left(h^{2}\right)\bm{\eta}_{\mathbf{x}}$
@@
(88) $=h\left(m_{1}\rho q-\frac{2m_{1}m_{2}\rho^{3}q^{2}}{\pi^{d/2}+2m_{2}\rho^{2}q}\right)\bm{\eta}_{\mathbf{x}}+h^{2}\widetilde{m}_{2}\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}+\mathcal{O}\left(h^{2}\right)\bm{\eta}_{\mathbf{x}}$
@@
(89) $=h\left(\frac{\pi^{d/2}m_{1}\rho q}{\pi^{d/2}+2m_{2}\rho^{2}q}\right)\bm{\eta}_{\mathbf{x}}+h^{2}\widetilde{m}_{2}\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}(\rho q)\bm{\eta}_{\mathbf{x}_{\perp}}+\mathcal{O}\left(h^{2}\right)\bm{\eta}_{\mathbf{x}}$
(90) $h\left(\frac{\pi^{d/2}m_{1}\rho q}{\pi^{d/2}+2m_{2}\rho^{2}q}\right)\bm{\eta}_{\mathbf{x}}+\frac{h^{2}\pi^{d/2}\rho}{2}\frac{\partial_{\bm{\eta}_{\mathbf{x}_{\perp}}}q}{1+\rho^{2}q}\bm{\eta}_{\mathbf{x}_{\perp}}+\mathcal{O}\left(h^{2}\right)\bm{\eta}_{\mathbf{x}}$
(91) $=\frac{h}{\rho}\left(\frac{\pi^{d/2}m_{1}\zeta}{\pi^{d/2}+2m_{2}\zeta}\right)\bm{\eta}_{\mathbf{x}}+\mathcal{O}\left(h^{2}\right)\bm{\eta}_{\mathbf{x}}+\mathcal{O}\left(h^{2}\right)\bm{\eta}_{\mathbf{x}_{\perp}}.$